# Nonlinear Propagation of Non-Gaussian Uncertainties: Theory and Applications to Spacecraft Trajectory Design

### v1

@@
Nonlinear Propagation of Non-Gaussian Uncertainties: Theory and Applications to Spacecraft Trajectory Design
@@ 1 Introduction
The role of uncertainty propagation is crucial in many fields of space trajectory design, ranging from low-Earth orbit determination, satellite tracking and collision avoidance, robust trajectory estimation in deep space missions, and space situational awareness [1]. Approaches to uncertainty propagation have ranged from trying to directly tackle the Fokker-Planck equation [2, 3], via approximated or numerical solutions, to methods leveraging just a first-order expansion of the dynamics (i.e., linear covariance propagation) [4], polynomial chaos expansion [5], and others related to statistical properties of Keplerian orbits [6, 7], or Gaussian mixture models combined with unscented transformation [8]. One of the most popular methods is via state transition tensors (STT), that leverage the high-order expansion of the flow around a nominal trajectory to map uncertainties at future times [9, 10]. However, all these STT methods have only dealt with initial Gaussian distributions, thereby preventing the treatment of generic probability density functions on both initial conditions and problem parameters.
While state transition tensors can be computed via the numerical integration of the variational equations (at high order), different approaches have proposed the use of alternative techniques to perform this task. These range from generalized dual numbers (used in the pyaudi library, which we extensively use in this work  [11]) to differential algebra [12, 13]. Using the variational approach has a different computational burden and numerical error associated, but is mathematically equivalent to these techniques as it produces the same high-order expansion (see Gimeno et al. [14]). Hence, to avoid further confusion, we will only use the term state transition tensor, referring to the quantities that represent the high-order derivatives of the flow with respect to the initial conditions and problem parameters.
In this work, we significantly extend the state transition tensor approach of Park & Scheeres [9] to non-Gaussian probability density functions (PDFs) for nonlinear propagation of uncertainties in dynamical systems with random initial conditions and parameters (e.g., initial position, gravitational constants, etc.). In the past, nonlinear uncertainty propagation has been mostly studied for distributions that are initially Gaussian or described by mixtures of Gaussian [15, 16, 17, 18, 19, 20]. Our approach allows dropping the Gaussian assumption, as generic mappings for any type of PDF can be found, provided that the PDF admits a moment generating function. Furthermore, we apply map-inversion techniques [21] to propagate uncertainties forward or backward in time at some predefined event, such as the crossing of a convenient or user-defined surface of section. We demonstrate that this approach enables robust trajectory design, whereby Gaussian and non-Gaussian uncertainties can be handled simultaneously in a variety of orbital mechanics problems.
@@ 2.1.1 Background
In summary, Eq. (3), which we also refer to as a high order Taylor map, provides a way to map initial perturbations around a nominal trajectory, at future times, provided that the expansion of the flow around the nominal trajectory has been computed. For $k\rightarrow\infty$, even if the Taylor series converges at the expansion point, its convergence will not be guaranteed at any distance from the nominal trajectory. The largest radius from the nominal trajectory at which the series still converges is referred to as the radius of convergence. While precise results exist for bounding the radius of convergence in Taylor series expansions, when considering a truncated series one needs to resort to approximated results [24]. This is discussed in the following section.
@@ 2.1.2 Radius of Convergence
where $b_{k}$ are the matrices resulting from unfolding the state transition tensor associated with the $k$-th Taylor term and ||.||||.||| | . | | indicates some norm, for example the 2-norm [28]. On the other hand, the Cauchy-Hadamard convergence radius, for multivariate cases, can be written as [29]:
@@ 2.1.3 Computing the Taylor Coefficients
Before delving into the connection between the high order Taylor maps introduced and the propagation of statistical moments, it is essential to clarify some practical details regarding the computation of this Taylor expansion. In literature, two main methods exist for the computation of the partial derivatives of the flow with respect to initial conditions and parameters: the application of some flavour of automatic differentiation techniques through a numerical integrator, which can be done via modern machine learning tools [30, 31] or via the algebra of truncated Taylor polynomials [32, 33, 11], or the direct numerical integration of the high-order variational equation [9, 22]. In [14], the authors show the mathematical equivalence of the variational equation approach to high-order automatic differentiation, for the same numerical integrator scheme. Hence, given a numerical integrator scheme, the same polynomial expansions should be obtained. While this is true mathematically, the implementation details can play a big difference in the computational time needed to compute these quantities practically. In previous work [34], for instance, it was found that the use of a differential algebraic approach (in that context, an automatic differentiation technique based on the algebra of truncated Taylor polynomials) is orders of magnitude more efficient than the variational one. In this work, instead, we opted for the variational approach and found that its efficiency is actually competitive with differential algebraic techniques. The contrast to what was found previously by Valli et al. [34] is to be attributed to the assumptions made in Valli’s work on the complexity of solving variational equations. There, such complexity was assumed to scale linearly with the number of equations to be integrated, an assumption which falls short in the case of the integration scheme employed in heyoka  [^1] https://bluescarni.github.io/heyoka/, date of access: May 2024. (the variational ODE solver used for this work) for the variational equations.
@@ 2.2 Probability Distributions
To compute the probability that random variable vectors take on a value in any given interval, one first needs to solve the above equation, and then integrate the PDF over that interval. However, the above equation is a partial differential equation that becomes particularly cumbersome to solve, especially for nonlinear dynamics and high-dimensional problems. Instead, we work with only the first moments of the distribution: these are generally descriptive of the PDF and do not require the full solution of the Fokker-Planck equation. In Sec. 2.2.2, we will introduce several definitions associated with these quantities.
@@ 2.2.3 Statistical Moments
$\begin{split}m_{3,ijk}&=\operatorname{\mathbb{E}}[(x_{i}-\operatorname{\mathbb{E}}[x_{i}])(x_{j}-\operatorname{\mathbb{E}}[x_{j}])(x_{k}-\operatorname{\mathbb{E}}[x_{k}])]\\ &=\operatorname{\mathbb{E}}[x_{i}x_{j}x_{k}]+2\operatorname{\mathbb{E}}[x_{i}]\operatorname{\mathbb{E}}[x_{j}]\operatorname{\mathbb{E}}[x_{k}]+\ -\operatorname{\mathbb{E}}[x_{i}]\operatorname{\mathbb{E}}[x_{j}x_{k}]-\operatorname{\mathbb{E}}[x_{j}]\operatorname{\mathbb{E}}[x_{i}x_{k}]-\operatorname{\mathbb{E}}[x_{k}]\operatorname{\mathbb{E}}[x_{i}x_{j}]\text{.}\end{split}$ (10)
@@ 2.3.1 General Formulation
$m_{2,ij}=\operatorname{\mathbb{E}}[(\delta x_{f}^{i}-\operatorname{\mathbb{E}}[\delta x_{f}^{i}])(\delta x_{f}^{j}-\operatorname{\mathbb{E}}[\delta x_{f}^{j}])]\approx\sum_{|\alpha|=1}^{k}\sum_{|\beta|=0}^{k}\dfrac{1}{\alpha!}\dfrac{1}{\beta!}(\partial^{\alpha}x_{f}^{i})\bigg{|}_{(\overline{\boldsymbol{x}}_{0},\overline{\boldsymbol{q}})}(\partial^{\beta}x_{f}^{j})\bigg{|}_{(\overline{\boldsymbol{x}}_{0},\overline{\boldsymbol{q}})}\operatorname{\mathbb{E}}[\delta\boldsymbol{z}^{\alpha}\delta\boldsymbol{z}^{\beta}]-\operatorname{\mathbb{E}}[\delta x_{f}^{i}]\operatorname{\mathbb{E}}[\delta x_{f}^{j}]\text{.}$ (21)
where $|\beta|=\sum_{s=0}^{M}\beta_{s}$ must also be taken over all possible combinations of $\beta_{s}\in\mathbb{N}$. Note that we take advantage of the expression derived in Eq. (9) to express the second moment, and that similar expressions can be derived for higher ones. These expressions will involve the computations of the higher order moment about the origin (see, for example, Eq. (10)), plus a series of coefficients that can all be computed from the lower-order moments.
@@
As a final note, it is important to also notice that with the proposed method, the $m$-th order moment will be composed of terms that are generated by the product of $m$ different polynomials, each expanded up to order $k$. This will produce, as a result, polynomials of increasing degree (i.e., by multiplying two polynomials of order $k$, one obtains a polynomial of order $k^{2}$), which will therefore have more and more terms that are negligible (since we are truncating at order $k$). This effect will necessarily cause, at a truncation fixed order $k$, a degradation of accuracy as higher moments are approximated.
@@ 2.3.2 Mapping Statistical Moments via High-order Maps
The formalism introduced above can be applied to map statistical moments of any given PDF through polynomial expressions, provided that the PDF admits a moment generating function. When using State Transition Tensors these polynomial expressions represent the flow of our dynamical system, but we can also obtain polynomials to represent the state of an arbitrary event equation (e.g. this could be the mapping between surfaces of section crossings in a dynamical system). The resulting Event Transition Tensors (ETTs) are computed as follows.
@@
Instead, having access to the Taylor expansion of the state at future times enables us to construct an accurate representation of the nonlinear event map by means of inversion techniques [44, 45]. Figure  1 illustrates two examples in which the moments are mapped either at an event represented by a surface of section (in this case a plane) of a periodic orbit (on the left), or for the case in which the event is constituted by the whole surface of a target asteroid body (on the right).
@@
Figure 1: Two schematic examples of statistical moments (first and second in this case) mapping onto the Poincaré map in a periodic orbit, on the left, and the case of the landing of a spacecraft onto the surface of an asteroid (on the right).
Without loss of generality, we consider the case in which the surface of the section is defined as the return map for one of the coordinates, meaning that it verifies $x_{f}^{j}-\overline{x}_{0}^{j}=0$, for a given $j$. Hence, by expanding the final state perturbation as done in Eq. (3), including the integration time as one of the parameters, then we can partially invert the map, and find the Taylor polynomial of the integration time corrections needed to ensure that any perturbation around the initial state and parameter returns to the surface of section:
@@
where $\delta t^{*}$ indicates the period correction that enables the return at the surface of section. Moreover, by substituting this Taylor polynomial into Eq. (3), one can obtain an expression for the reduced state (i.e., the $n-1$ state coordinates at the surface of section): this will also be a Taylor polynomial of order $k$. Hence, at the only additional cost of partially inverting the map and making a few substitutions, and using the same approach detailed in Sec. 2.3.1, one can directly propagate the statistical moments at the event manifold.
@@ 3.1.1 Unperturbed Case
We compare the statistics reconstructed via the moment generating function method we propose, and the sample statistics reconstructed via Monte Carlo propagation. In particular, we compute the Monte Carlo statistics by sampling 10 million points from the uncertain PDF of the state and parameter and propagating each of them at the final integration time. This is compared against the method described in Sec. 2.3.1, which only requires a single Taylor expansion around the nominal state, combined with the cost of computing the expectation over the initial probability density functions (which, however, can be performed only once, offline and independently of the specific dynamics). While depending on the chosen order of the expansion, the computational cost can increase substantially: in our case, we find that for all the tested orders (i.e., up to order 4), the polynomial map can be very quickly constructed, at orders of magnitude less computation time than the Monte Carlo approach.
In Fig. 2, we display 15 projections of the final reconstructed statistics, in terms of mean and covariance, using different Taylor order expansions of our proposed method, against the ones from the Monte Carlo method. Moreover, in the lower right plot, we also display (in orange) the three-dimensional Monte Carlo-sampled states as they are propagated along the orbit. As it can be seen, for all projections, from order 2 and above, the technique manages to very accurately capture both the mean and covariance matrix of the underlying probability density function. Furthermore, as we observe from the orange and blue samples, which represent the Monte Carlo trajectories before and after the integration time, respectively, the probability density function does not start from a Gaussian, nor converges towards it after the integration. To diagnose the non-normality of the propagated distribution, one can use moments above the second: for instance, the third moment provides information about the skewness of the distribution. In Fig. 3, we display the accuracy as a function of integration times at which the proposed technique can reconstruct the first three moments at a 4th-order Taylor expansion (on the left), together with the accuracy at which the second statistical moment (i.e., the covariance) is reconstructed at different orders, ranging from the first to the 4th (on the right). The error in the reconstruction of the $k$-th statistical moment is computed via the absolute error:
@@
In both Eq. (24) and (25), we use $m_{k,t}$ to indicate the true $k$-th statistical moment tensor, which in this case is computed via Monte Carlo, and $m_{k,est}$ to indicate the approximated $k$-th statistical moment tensor, which in this case is computed by the method introduced in Sec. 2.3.1. For order one (i.e., $k=1$), the norm is a standard 2-norm of a vector, however, for higher order, this becomes the Frobenius norm (due to the fact that the moments become tensors). As observed from Fig. 3, the method effectively captures the first three moments with relative errors below 1% in the reconstruction of the moment tensors. The relative error is useful as it provides a direct interpretation in terms of percentage, allowing for an intuitive assessment of the reconstruction accuracy. However, the utility of relative error diminishes when the denominator, i.e., the norm of the $k$-th order moment, approaches zero. In such cases, the relative error can become disproportionately large or even undefined, rendering it meaningless. To address this limitation, one can resort to the absolute error, which we also display in the top row of the figure. This remains meaningful in scenarios where the moment norm goes to zero.
Figure 2: Moments statistics and projections of the six-dimensional state components propagated via Monte Carlo and performed with the proposed method, at different Taylor expansion orders.
Figure 3: Left: error (absolute on top, relative bottom) of reconstruction (between the proposed method and Monte Carlo) in the first three statistical moments as a function of propagation time; right: error (absolute on top, relative bottom) of reconstruction of the second moment, for different orders, as a function of time.

### v2

@@
Nonlinear Propagation of Non-Gaussian Uncertainties
@@ 1 Introduction
Uncertainty propagation plays a crucial role in various aspects of space trajectory design, including low-Earth orbit determination, satellite tracking, collision avoidance, robust trajectory estimation in deep space missions, and space situational awareness [1]. Approaches to uncertainty propagation vary widely, from attempts to solve the Fokker-Planck equation directly [2, 3], using approximated or numerical methods, to techniques based on first-order dynamics expansions (e.g., linear covariance propagation)[4], polynomial chaos expansion[5], and change of variable technique [6]. Other methods leverage statistical properties of Keplerian orbits [7, 8] or apply Gaussian mixture models combined with unscented transformations [9].
One of the most popular methods is via state transition tensors (STT), which leverages the high-order expansion of the flow around a nominal trajectory to map uncertainties at future times [10, 11]. However, methods based on STTs have so far only dealt with initial Gaussian distributions, thereby preventing the treatment of generic probability density functions on both initial conditions and problem parameters.
While STTs can be computed via the numerical integration of the variational equations, different approaches have proposed the use of alternative techniques to perform this task. These range from generalized dual numbers [12] to differential algebra [13, 14] to machine learning pipelines as popularized, for example, by the recent NeuralODE paradigm [15, 16]. According to the context, STTs are often referred to using a variety of different names and notations. We will only use the term state transition tensors, referring to the quantities that represent the high-order derivatives of the flow with respect to the initial conditions and problem parameters. With reference to the work of Park & Scheeres [10], we extend the ideas on uncertainty propagation through STTs to the case of non-Gaussian probability density functions (PDFs) associated to the dynamical system initial conditions and parameters (e.g., initial position, gravitational constants, etc.).
In the past, nonlinear uncertainty propagation has been primarily studied in the case of distributions that are initially either Gaussian or described by mixtures of Gaussian [17, 18, 19, 20, 21, 22]. Our approach allows dropping the Gaussian assumption, as generic mappings for any type of PDF can be found, provided that the PDF admits a moment-generating function. By using this approach to propagate uncertainties through STTs, we effectively also mitigate potential issues arising from distribution tails extending beyond the radius of convergence of the expanded flow, as we ensure that probability density functions have a compact domain fully contained within this radius. We test the proposed method leveraging map-inversion techniques [23] and propagating uncertainties forward or backward at predefined events, such as crossing a surface of section or similar.
@@ 2.1.1 Background
In summary, Eq. (3), which we also refer to as a high-order Taylor map, provides a way to map initial perturbations around a nominal trajectory, at future times.
The largest norm of the perturbation $\boldsymbol{\delta z}$ within which the series converges is referred to as the radius of convergence. When dealing with expansions such as that in Eq. (3), it is important to estimate the radius of convergence as well as the truncation error. We use an estimate of the radius of convergence to check that all uncertainties are within the convergence radius (a necessary condition that allows the use of the expansions for their propagation). As for the truncation error, we compute it exactly and report its value to cross-validate the specific computations presented here. Note that, in principle, the truncation error can also be bounded rigorously, e.g. using Taylor models [26], but not necessarily in a sufficiently tight enclosure.
@@ 2.1.2 Radius of Convergence
where $b_{k}$ are the matrices resulting from unfolding the STT associated with the $k$-th Taylor term and ||.||||.||| | . | | indicates some norm, for example the 2-norm [30]. On the other hand, the Cauchy-Hadamard convergence radius, for multivariate cases, can be written as [31]:
@@ 2.1.3 Computing the Taylor Coefficients
Before delving into the connection between the high order Taylor maps introduced and the propagation of statistical moments, it is essential to clarify some practical details regarding the computation of this Taylor expansion. In literature, two main methods exist for the computation of the partial derivatives of the flow with respect to initial conditions and parameters: the application of some flavor of automatic differentiation techniques through a numerical integrator, which can be done via modern machine learning tools [32, 16] or via the algebra of truncated Taylor polynomials [33, 34, 12], or the direct numerical integration of the high-order variational equation [10, 24]. In [35], the authors show the mathematical equivalence of the variational equation approach to high-order automatic differentiation, for the same numerical integrator scheme. Hence, given a numerical integrator scheme, the same polynomial expansions will be obtained. While this is true mathematically, the implementation details can play a big difference in the computational time needed to compute these quantities practically. In previous work [36], for instance, the use of differential algebra (in that context a forward mode automated differentiation technique) to compute the STTs was reported to be orders of magnitude more efficient than the variational approach. In contrast to that finding, we use the variational approach to compute STTs for all the cases presented in Sec. 3 and found that it is as competitive as differential algebraic techniques while allowing for the seamless use of Taylor integrators and the resulting reliable event detection machinery described in Ref. [37, 38]. This finding sharply contrasts with the conclusions drawn by Valli et al.[36], likely due to the assumption they made on the number of variational equations needed to compute the needed tensors. In Valli’s work, this is presumed to grow linearly with $N=\sum_{q=1}^{k+1}n^{q}$ where $n$ is the state dimension and $k$ is the expansion order. The same assumption was also shared by earlier studies, such as [10]. However, this scaling fails to apply to the variational equations we derive automatically using the heyoka variational ODE tool [^1] https://bluescarni.github.io/heyoka/, accessed October 2024. [37]. In this case, the number of variational equations $N$ incorporates all symmetries emerging from Schwarz’s theorem, thereby resulting in a significantly reduced count [39]:
$N=n\sum_{q=1}^{k}\left(\begin{array}[]{c}q+n-1\\ q\end{array}\right).$
The final number is orders of magnitude smaller than previously assumed as the order increases. For example, in the case of $n=6$ equations, varying $k$ up to the fifth order we have [42, 168, 504, 1260, 2772] variational equations, while previous assumptions expected this number to grow as [42, 258, 1554, 9330, 55986].
@@ 2.2 Probability Distributions
To compute the probability that random variable vectors take on a value in any given interval, one first needs to solve the above equation, and then integrate the PDF over that interval. While simplifications arise when the dynamical system is Hamiltonian because, according to Liouville’s theorem, the phase-space probability density function remains constant along the system’s trajectories, the above equation is a partial differential equation that becomes particularly cumbersome to solve, especially for nonlinear dynamics and high-dimensional problems. Instead, we work with only the first moments of the distribution: these are generally descriptive of the PDF and do not require the full solution of the Fokker-Planck equation. In the next section, we will introduce several definitions associated with these quantities.
@@ 2.2.3 Statistical Moments
$\begin{split}m_{3,ijk}&=\operatorname{\mathbb{E}}[(x_{i}-\operatorname{\mathbb{E}}[x_{i}])(x_{j}-\operatorname{\mathbb{E}}[x_{j}])(x_{k}-\operatorname{\mathbb{E}}[x_{k}])]\\ &=\operatorname{\mathbb{E}}[x_{i}x_{j}x_{k}]+2\operatorname{\mathbb{E}}[x_{i}]\operatorname{\mathbb{E}}[x_{j}]\operatorname{\mathbb{E}}[x_{k}]+\\ &\ \ -\operatorname{\mathbb{E}}[x_{i}]\operatorname{\mathbb{E}}[x_{j}x_{k}]-\operatorname{\mathbb{E}}[x_{j}]\operatorname{\mathbb{E}}[x_{i}x_{k}]-\operatorname{\mathbb{E}}[x_{k}]\operatorname{\mathbb{E}}[x_{i}x_{j}]\text{.}\end{split}$ (10)
@@ 2.3.1 General Formulation
$\begin{split}m_{2,ij}&=\operatorname{\mathbb{E}}[(\delta x_{f}^{i}-\operatorname{\mathbb{E}}[\delta x_{f}^{i}])(\delta x_{f}^{j}-\operatorname{\mathbb{E}}[\delta x_{f}^{j}])]\approx\\ &\sum_{|\alpha|=1}^{k}\sum_{|\beta|=0}^{k}\dfrac{1}{\alpha!}\dfrac{1}{\beta!}(\partial^{\alpha}x_{f}^{i})\bigg{|}_{(\overline{\boldsymbol{x}}_{0},\overline{\boldsymbol{q}})}(\partial^{\beta}x_{f}^{j})\bigg{|}_{(\overline{\boldsymbol{x}}_{0},\overline{\boldsymbol{q}})}\operatorname{\mathbb{E}}[\delta\boldsymbol{z}^{\alpha}\delta\boldsymbol{z}^{\beta}]+\\ &-\operatorname{\mathbb{E}}[\delta x_{f}^{i}]\operatorname{\mathbb{E}}[\delta x_{f}^{j}]\end{split}$ (21)
where $|\beta|=\sum_{s=0}^{M}\beta_{s}$ must also be taken over all possible combinations of $\beta_{s}\in\mathbb{N}$. Note that we take advantage of the expression derived in Eq. (9) to express the second moment, and that similar expressions can be derived for higher ones. These expressions will involve the computations of the higher-order moment about the origin (see, for example, Eq. (10)), plus a series of coefficients that can all be computed from the lower-order moments.
@@
As a final note, we highlight that, with the proposed method, the $m$-th order moment will be composed of terms that are generated by the product of $m$ different polynomials, each expanded up to order $k$. This will produce, as a result, polynomials of increasing degree (i.e., by multiplying two polynomials of order $k$, one obtains a polynomial of order $k^{2}$). If one were to work with the algebra of truncated polynomials, then truncating at order $k$ would cause all higher-order contributions to disappear, which would necessarily cause a degradation of accuracy. Hence, in these cases, it is important to retain higher-order terms.
@@ 2.3.2 Mapping Statistical Moments via High-order Maps
The formalism introduced above can be applied to map statistical moments of any given PDF through polynomial expressions, provided that the PDF admits a moment generating function. When using STTs these polynomial expressions represent the flow of our dynamical system, but we can also obtain polynomials to represent the state of an arbitrary event equation (e.g. this could be the mapping between surfaces of section crossings in a dynamical system). The resulting Event Transition Tensors (ETTs) are computed as follows.
@@
Instead, having access to the Taylor expansion of the state at future times enables us to construct an accurate representation of the nonlinear event map by means of inversion techniques [49, 50, 14]. Figure  1 illustrates two examples in which the moments are mapped either at an event represented by a surface of section (in this case a plane) of a periodic orbit (on the left), or for the case in which the event is constituted by the whole surface of a target asteroid body (on the right).
@@
Figure 1: Schematic examples of mapping of moments onto a Poincaré map (left) and onto the surface of an asteroid (right).
Without loss of generality, we consider the case in which the surface of the section is defined as the return map for one of the coordinates, meaning that it verifies $x_{f}^{j}-\overline{x}_{0}^{j}=0$, for a given $j$. As already discussed in previous work (e.g.,  [51] and references therein), by expanding the final state perturbation as done in Eq. (3) and including the integration time as one of the parameters, the map can be partially inverted, and the Taylor polynomial of the integration time corrections needed to ensure that any perturbation around the initial state and parameter returns to the surface of section can be found:
@@
where $\delta t^{*}$ indicates the period correction that enables the return at the surface of section. Moreover, by substituting this Taylor polynomial into Eq. (3), one can obtain an expression for the reduced state (i.e., the $n-1$ state coordinates at the surface of section): this will also be a Taylor polynomial of order $k$. Hence, at the only additional cost of partially inverting the map and making a few substitutions, and using the same approach detailed above, one can directly propagate the statistical moments at the event manifold.
@@ 3.1.1 Unperturbed Case
We compare the statistics reconstructed via the moment generating function method we propose, and the sample statistics reconstructed via Monte Carlo propagation. In particular, we compute the Monte Carlo statistics by sampling 10 million points from the uncertain PDF of the state and parameter and propagating each of them at the final integration time. This is compared against the method described in Sec. 2.3, which only requires a single Taylor expansion around the nominal state, combined with the cost of computing the expectation over the initial probability density functions (which, however, can be performed only once, offline and independently of the specific dynamics). While depending on the chosen order of the expansion, the computational cost can increase substantially: in our case, we find that for all the tested orders (i.e., up to order 4), the polynomial map can be very quickly constructed, at orders of magnitude less computation time than the Monte Carlo approach.
In Fig. 2, we display 4 projections of the final reconstructed statistics, in terms of mean and covariance, using different Taylor order expansions of our proposed method, against the ones from the Monte Carlo method. In all figures, the order 1 estimation of the covariance is off, while orders 2 and 3 overlap and almost converge towards the ground truth distribution (the Monte Carlo one, in this case), and order 4 overlaps with the ground truth. Moreover, in the lower right plot, we also display (in orange) the three-dimensional Monte Carlo-sampled states as they are propagated along the orbit. As it can be seen, for all projections, from order 2 and above, the technique manages to very accurately capture both the mean and covariance matrix of the underlying probability density function. Furthermore, as we observe from the orange and blue samples, which represent the Monte Carlo trajectories before and after the integration time, respectively, the probability density function does not start from a Gaussian, nor converges towards it after the integration. To diagnose the non-normality of the propagated distribution, one can use moments above the second: for instance, the third moment provides information about the skewness of the distribution. In Fig. 3, we display the accuracy as a function of integration times at which the proposed technique can reconstruct the first three moments at a 4th-order Taylor expansion (on the left), together with the accuracy at which the second statistical moment (i.e., the covariance) is reconstructed at different orders, ranging from the first to the 4th (on the right). The error in the reconstruction of the $k$-th statistical moment is computed via the absolute error:
@@
In both Eq. (24) and (25), we use $m_{k,t}$ to indicate the true $k$-th statistical moment tensor, which in this case is computed via Monte Carlo, and $m_{k,est}$ to indicate the approximated $k$-th statistical moment tensor, which in this case is computed by the method introduced in Sec. 2.3. For order one (i.e., $k=1$), the norm is a standard 2-norm of a vector, however, for higher order, this becomes the Frobenius norm (due to the fact that the moments become tensors). As observed from Fig. 3, the method effectively captures the first three moments with relative errors below 1% in the reconstruction of the moment tensors. The relative error is useful as it provides a direct interpretation in terms of percentage, allowing for an intuitive assessment of the reconstruction accuracy. However, the utility of relative error diminishes when the denominator, i.e., the norm of the $k$-th order moment, approaches zero. In such cases, the relative error can become disproportionately large or even undefined, rendering it meaningless. To address this limitation, one can resort to the absolute error, which we also display in the top row of the figure. This remains meaningful in scenarios where the moment norm goes to zero.
Figure 2: Monte Carlo and Taylor expansion approximation of the moments for the two body unperturbed cases.
Figure 3: Errors as a function of time in the first three moments (left), and of the second moment for different orders (right).